# Autoregressive Ranking: Bridging the Gap Between Dual and Cross Encoders

### v1

@@ 3.2 Ranking via (Pointwise) Generative Rankers
where $\operatorname{E}_{\text{docIDs}}\in\mathbb{R}^{|\mathscr{V}_{\text{docIDs}}|\times n}$ represents a matrix whose rows are the token embeddings associated with $\mathscr{V}_{\text{docIDs}}$. To study the rank requirements of $\operatorname{E}_{\text{docIDs}}$, we will introduce for convenience the matrix $\operatorname{E}^{\prime}\coloneqq[\operatorname{E}_{\text{docIDs}}\mathbf{1}_{|\mathscr{V}_{\text{docIDs}}|}]\in\mathbb{R}^{|\mathscr{V}_{\text{docIDs}}|\times(n+1)}$, i.e., $\operatorname{E}_{\text{docIDs}}$ augmented with a column of ones.
@@
To investigate the representational capacity of this model, we consider an idealized “infinite-capacity” setting where the hidden vector $\varphi(c)$ can be any vector in $\mathbb{R}^{n}$, which implies the following:
Proposition 3.1.
@@
An infinite-capacity ARR can produce any (strictly positive) probability distribution over tokens in $\mathscr{V}_{\text{docIDs}}$ via constrained decoding if and only if $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
@@
The contrapositive of the forward implication ($\Rightarrow$) is commonly known as the “softmax bottleneck” (yang2017breaking):
@@
Recall from Equation˜6 that $\sum_{i\in\mathcal{P}}h_{i}z^{\prime}_{i}=\sum_{i\in\mathcal{N}}|h_{i}|z^{\prime}_{i}$. Now,
@@
Since $aS<bS$, we get $\sum_{i\in\mathcal{P}}h_{i}z^{\prime}_{i}<\sum_{i\in\mathcal{N}}|h_{i}|z^{\prime}_{i}$, which is a contradiction ($\bot$) with Equation˜6.
@@
In particular, we note that many existing language models come with “special tokens” which can be used as $\mathscr{V}_{\text{docIDs}}$, and for any existing language model which does not come with pre-allocated “special tokens” we can expand the vocabulary with $n$ new tokens and extend its embedding matrix with the $I_{|\mathscr{V}_{\text{docIDs}}|}$ identity matrix, which would satisfy the requirements. An implicit limitation of this is that $|\mathscr{V}_{\text{docIDs}}|\leq n$, however the chain-rule of probability implies the following.
Corollary 3.1.
An infinite-capacity ARR can produce any (strictly positive) probability distribution over sequences of tokens in $\mathscr{V}_{\text{docIDs}}$ via constrained decoding if and only if $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
The capacity to produce arbitrary probability distributions over sequences clearly entails the ability to produce arbitrary rankings over sequences. One may wonder, however, whether the requirement of ˜3.1/Corollary˜3.1 is also necessary for ranking. Surprisingly, at a single token position the requirement for ranking is the same as for probability distributions:
Proposition 3.2.
An infinite-capacity ARR can produce any permutation over tokens in $\mathscr{V}_{\text{docIDs}}$ via constrained decoding if and only if $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
@@
($\Rightarrow$) The proof is entirely analogous to the forward proof of ˜3.1, with the exception that the contradiction targets permutations in which all token indices for which $h$ takes negative values precede all token indices for which $h$ takes negative values.

### v2

@@ 3.2 Ranking via (Pointwise) Autoregressive Rankers
where $\operatorname{E}_{\text{docIDs}}\in\mathbb{R}^{|\mathscr{V}_{\text{docIDs}}|\times n}$ represents a matrix whose rows are the token embeddings associated with $\mathscr{V}_{\text{docIDs}}$. To study the rank requirements of $\operatorname{E}_{\text{docIDs}}$, we introduce for convenience the matrix $\operatorname{E}^{\prime}\coloneqq[\operatorname{E}_{\text{docIDs}}\mathbf{1}_{|\mathscr{V}_{\text{docIDs}}|}]\in\mathbb{R}^{|\mathscr{V}_{\text{docIDs}}|\times(n+1)}$, i.e., $\operatorname{E}_{\text{docIDs}}$ augmented with a column of ones.
@@
We consider an idealized “infinite-capacity” setting where the hidden vector $\varphi(c)$ can be any vector in $\mathbb{R}^{n}$.
{restatable}
@@
[Arbitrary Distributions over Tokens]propPropArbitraryTokenLevelProb An infinite-capacity ARR can produce any (strictly positive) probability distribution over tokens in $\mathscr{V}_{\text{docIDs}}$ via constrained decoding iff $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
@@
The contrapositive of the forward implication is commonly known as the “softmax bottleneck” (yang2017breaking).
We note that many existing language models come with “special tokens” which can be used as $\mathscr{V}_{\text{docIDs}}$, and for any existing language model which does not come with pre-allocated “special tokens” we can expand the vocabulary with $n$ new tokens and extend its embedding matrix with the $I_{|\mathscr{V}_{\text{docIDs}}|}$ identity matrix, which would satisfy the requirements. An implicit limitation of this is that $|\mathscr{V}_{\text{docIDs}}|\leq n$, however the chain-rule of probability implies the following.
Corollary 3.1 (Arbitrary Distributions over Sequences).
An infinite-capacity ARR can produce any (strictly positive) probability distribution over sequences of tokens $\mathscr{V}_{\text{docIDs}}$ via constrained decoding iff $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
Producing arbitrary probability distributions over sequences clearly entails the ability to produce arbitrary rankings over sequences. One may wonder, however, whether the requirement of Section˜3.2 / Corollary˜3.1 is also necessary for ranking. Surprisingly, at a single token position the requirement for ranking is the same as for distributions:
{restatable}
[Arbitrary Permutations over Tokens]propPropArbitraryTokenLevelRanking An infinite-capacity ARR can produce any permutation over tokens in $\mathscr{V}_{\text{docIDs}}$ via constrained decoding iff $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$.
We observe that the capacity to produce arbitrary permutations over sequences depends on the capacity to express all $|\mathscr{V}_{\text{docIDs}}|!$ rankings at a single token position. Since the requirement for this is the same as the requirement to represent arbitrary distributions at both token- and sequence-level, we conclude that the rank requirement expressed in Section˜3.2, Corollary˜3.1, and Corollary˜3.1 is not only sufficient but necessary for the ranking task.
Remark: Is the theoretical requirement on $\operatorname{E}^{\prime}$ satisfied by LLMs used in practice for reranking, and why are those LLMs successful even when this requirement clearly can not be satisfied? Our experiments on WordNet and ESCI (Section˜5.1) reflect two common scenarios. In WordNet, where docIDs comprise English nouns, the docID tokens span almost all of $|\mathcal{V}|$. Since the model dimension is much smaller, $\operatorname{E}^{\prime}$ can not achieve a rank even close to $|\mathscr{V}_{\text{docIDs}}|$. [^4] Mistral-7B-v0.3-it has $|\mathcal{V}|=32,768$ and $d=4,096$. However, most of the possible sequences are simply never encountered, and the model is tasked with learning a very small subset of all possible permutations. In ESCI, by contrast, we choose numerical docIDs, and the resulting $|\mathscr{V}_{\text{docIDs}}|$ is therefore closer to the model dimension, allowing the rank requirement on $\operatorname{E}^{\prime}$ to be matched more accurately.
At either the single token position or the sequence level, we see that the same requirement on the docID token embeddings (namely, that $\operatorname{rank}(\operatorname{E}^{\prime})=|\mathscr{V}_{\text{docIDs}}|$, i.e., slightly weaker than $\operatorname{E}_{\text{docIDs}}$ being full-rank) is both necessary and sufficient to enable a fixed-dimension ARR to generate arbitrary probability distributions or arbitrary permutations for an unrestricted number of documents. This ability is strictly superior to DEs, whose embedding dimension must grow linearly with the number of documents to achieve arbitrary rankings.
4 Generalized Rank-Aware Training Loss
Given an example “point” of the form $(q,d_{r},r)$ with the query, a docID, and the true rank, we propose a generalized loss function for autoregressive ranking which can be viewed as a weighted item-level loss, with the weight coefficient $\lambda(r)$ for a given docID being a function of the document’s rank $r$. The item-level loss decomposes as a sum of cross entropy losses between the model predictions and target distributions y at each timestep (token position) of the docID under consideration. Note that the target distributions are a function of the timestep $t$ and potentially also of the current docID’s rank $r$, hence $\mathbf{y}(r,t)$. Our generalized next-token prediciton loss for ranking, SToICaL (for “Simple Token-Item Calibrated Loss”), is given below. $\theta$ represents model parameters, $\mathcal{T}$ the tokenizer, and $\bar{q}$ the “augmented query” (cf. Section˜5.1):
The most common target distribution $\mathbf{y}(r,t)$ is one-hot:
Section˜4.2 presents a rank-aware improvement to $\mathbf{y}(r,t)$.
4.1 Rank-Aware Item-Level Reweighting
4.2 Prefix Tree for Rank-Aware Token-Level $\mathbf{y}(r,t)$
A prefix tree (aka trie) groups strings according to their shared prefixes for quickly determining (by traversing left-to-right, i.e., from the root down) whether a given sequence belongs to the valid set of sequences or not. Building a prefix tree over the list of tokenized docIDs allows us to examine, at each timestep, which continuation docIDs are permissible given the partially-generated docID, and to distribute mass accordingly among valid tokens in the target distribution.
Given tokenized target docIDs in the following toy example:
[…]
we can construct a prefix tree over their tokens, as shown in Figure˜2(a). The leaf nodes and potentially some nonterminal nodes of the trie, which we term docID nodes and represent with solid borders in Figure˜2(a), will each represent the tokenization of a docID $d_{r}$, and to such nodes we assign the score $\mu(r)=\frac{1}{r^{\beta}}$ for temperature hyperparameter $\beta$. To score any given node in the trie (e.g., for a partially-generated docID), we can marginalize over the scores of all docIDs which that node governs. [^5] At marginalization, the fractional $\mu(r)$ prevents a continuation which leads to the top-$r$ docID from being outweighed by a continuation leading to multiple lower-ranked docIDs. As $\beta\rightarrow\infty$, the resulting distributions converge to one-hots.
Figure 2: Illustration of (a) a prefix-tree; target distributions derived from it via (b) one-hot supervision versus (c) marginalization.
Given an empty string ($\varepsilon$) at timestep $t_{1}$, the trie tells us that permissible continuation tokens for $\varepsilon$] are $\{\texttt{c, d, f}\}$. At timestep $t_{2}$, if our highest-ranked document is $d_{1}=\texttt{``dog''}$, we will have teacher-forced token d as the $t_{2}$ input to the model, and the prefix tree will consequently yield tokens $\{\texttt{og, eer}\}$ as valid continuations for $\varepsilon$, d]. We illustrate the (pre-normalized) target distributions produced by this trie-based marginalization process in Figure˜2(c). [^6] Technically, for timesteps $t>1$, teacher forcing causes a mismatch between the rank-aware training described here and the model’s actual decision-making at inference time. For example, if the model chooses ‘‘c’’ at timestep t=1 during inference, we will not have smoothed over ‘‘at’’ at $t=2$ during training to encourage the second-best ‘‘cat’’ to be generated (without constrained decoding, we might well generate ‘‘c’’ ++ ‘‘og’’ $\rightarrow$ ‘‘cog’’ in such a scenario). Nevertheless, akin to reddy2024first, our method does encourage full rank-awareness at $t=1$. We leave the question of mitigating this rank-aware training-inference mismatch for $t>1$ to future work.
5 Experiments
5.1 Datasets, Experimental Settings, and Metrics
WordNet
  • Query: deer
  • Ranked docIDs: ruminant $\succ$ even-toed ungulate $\succ$ ungulate $\succ$ placental $\succ$ mammal $\succ$ vertebrate $\succ$ chordate $\succ$ animal $\succ$ organism $\succ$ living thing $\succ$ whole $\succ$ object $\succ$ physical entity $\succ$ entity]
  • Negative docID: solarization
We provide a zero-shot prompt for WordNet in Section˜D.1.
ESCI Shopping Queries
Unlike the WordNet dataset (Section˜5.1), the documents of ESCI are (a) substantially longer, i.e., on the order of dozens of tokens instead of ten, and (b) and have not been sampled from a taxonomy. Issue (a) motivates the creation of docIDs that are both short enough to generate as pointwise targets, and structured enough so that a prefix tree can be built over them for target training distributions as discussed in Section˜4.2. We create such docIDs via sparse dictionary learning over the Gecko product title embeddings to obtain sparse vectors, each of dimension 100 with 3 nonzero values. [^7] We use Orthogonal Matching Pursuit from scikit-learn (pedregosa2011scikit), n_components=100 and transform_n_nonzero_coefs=3. The docID for a given product title is then taken to be the concatenation of the indices of the nonzero entries in its corresponding sparse vector, sorted in descending order by the absolute values of the entries (e.g., “25,36,39”). An zero-shot prompt for ESCI is provided in Section˜D.2.
Task Setup for Autoregressive Ranking
The in-context input to our Causal Transformer is a query $q$ and an unordered docID set $\operatorname{shuffle}(\{d_{1},\ldots,d_{n_{q}}\})$; assume the ground truth ranking is $d_{1}\succ\ldots\succ d_{n_{q}}$. We construct the prompt, denoted by $\bar{q}$ for the query augmented with the corpus and system prompt, as $\bar{q}\coloneqq\operatorname{prompt}(q,\operatorname{shuffle}(\{d_{1},\ldots,d_{n_{q}}\}))$, cf. Appendix˜D. We emphasize that this augmented prompt construction reflects an in-context approach to ranking (gupta2025scalable) in which the full corpus is provided as context — while this contrasts with DEs’ and CEs’ pairwise scoring approach, it is natural to ARR, which aims to exploit the in-context copying and reasoning abilities of decoder LLMs. For our evaluation, as a simple proxy to beam search, we perform greedy decoding for $\max_{i\in[1,n_{q}]}|\mathcal{T}(d_{i})|$ steps, where $\mathcal{T}$ is the tokenizer, and score any document by averaging the log-probabilities of its constituent tokens, i.e., $\operatorname{score}(d_{j})=\frac{1}{|\mathcal{T}(d_{j})|}\sum_{k=1}^{|\mathcal{T}(d_{j})|}\log p_{\theta}(\mathcal{T}(d_{j})[k]\mid\mathcal{T}(\bar{q}),\mathcal{T}(d_{j})_{<k})$. [^8] We emphasize that in practice, a $\operatorname{top}-k$ ranking can still be directly generated via beam search using beam size $k$. We use Mistral-7B-v0.3-it (jiang2024identifying) for all fine-tuning experiments in this paper, with settings for the generative approach in Section˜B.1.
$\alpha=2$ 0.0% 99.83 97.74 95.69 95.48 96.89 96.58
$\alpha=3$ 0.0% 99.81 99.22 97.35 96.16 96.07 95.53
$\alpha=4$ 0.02% 99.78 99.36 97.46 95.63 96.03 95.31
$\alpha=5$ 0.04% 99.67 99.6 97.14 94.55 95.71 93.52
$\lambda(r)=\frac{n_{q}-r+1}{n_{q}}$, $\mathbf{y}$ one-hot 0.0% 99.6 51.62 69.88 82.09 89.6 93.34
$\beta=2$ 1.38% 96.54 99.3 67.31 57.51 56.98 62.45
$\beta=3$ 2.04% 96.34 99.44 67.58 59.95 58.58 63.17
$\beta=4$ 4.54% 96.36 99.58 68.53 62.25 62.12 67.41
$\beta=5$ 13.04% 96.47 99.78 70.84 65.12 64.81 70.43
$\beta=2$ 97.21 70.32 58.06 51.07 48.08 54.96 69.03
$\beta=3$ 97.11 68.71 56.13 48.03 45.78 52.31 67.63
$\beta=4$ 96.96 69.03 54.84 44.6 43.9 50.29 67.7
$\beta=5$ 96.89 66.77 56.13 44.21 42.4 50.72 67.61
Metrics
  • We measure the normalized discounted cumulative gain (nDCG) between “gold” scores and predicted log-probabilities (including for the sampled negative document). The “gold” are in log-scale in decreasing order of relevance: $\operatorname{score}(d_{1}^{+})=\log(n_{q}+1),\operatorname{score}(d_{2}^{+})=\log(n_{q})\ldots,\operatorname{score}(d_{n_{q}-1}^{+})=\log(3),\operatorname{score}(d_{n_{q}}^{+})=\log(2),\operatorname{score}(d_{q}^{-})=\log(1)=0$.
  • We use the log-probabilities to $\operatorname{argsort}$ over the docIDs, and use the gold ranking $\pi$ and predicted ranking $\hat{\pi}$ to measure recall-at-k (R@k) as $\frac{|\pi[:k]\cap\hat{\pi}[:k]|}{k}$.
5.2 Main Results
In Tables 2 and 2 we present our experimental results on the WordNet and ESCI datasets, respectively. Each of our ranking-aware loss functions can be seen to drastically reduce the constraint violation rate (cf. Table˜2) — i.e., the model learns to score irrelevant docIDs lower than relevant docIDs for a query. Such a shift in probability mass from invalid to valid docIDs makes the resulting LLM more aligned, as it were “by construction”, with the constraints of generating valid docIDs. Such realignment would naturally ease the computational burden of a debiasing algorithm for constrained decoding, e.g., as described in ye2025efficient.
Fixing our y to one-hots, we observe for WordNet in Table˜2 that fractional ($\lambda(r)=\frac{1}{r^{\alpha}}$) item-level reweighting is more effective than stepwise ($\lambda(r)=\frac{n_{q}-r+1}{n_{q}}$), while in the limit $\alpha\rightarrow\infty$ for fractional reweighting, the target distributions look more and more like a top-1 training curriculum $\frac{1}{r^{\alpha}}\rightarrow\mathbb{I}_{r=1}$ (cf. Section˜4.1). We separately explore the effect of rank-aware target distributions $\mathbf{y}(r,t)$ (cf. Section˜4.2), with a single rank-aware target per query (rather than per document $d_{r}$). We similarly observe decreased constraint violation rate, improved nDCG and R@K for $K>2$, but not as dramatic as for the item-level reweighting loss.
Figure 3: Results on WordNet with Dual Encoders (DEs) and Cross Encoders (CEs). (a) Comparison of autoregressive ranking with DEs and CE. (b, c) Effect of $\alpha$ in the reweighting function for DEs and CEs, respectively.
The ESCI dataset contains a much higher number of product titles to be reranked per query on average than the WordNet dataset (cf. Appendix˜C). Training with item-level losses until convergence would treat each docID $d_{r}$ for $r\in[1,n_{q}]$ as a separate target, and would thus require a proportionally larger number of training steps to iterate over the entire dataset. Hence, we run our ESCI experiments under the same budget constraint as next-token prediction, providing a single target per every query by setting $\lambda(r)=\mathbb{I}_{r=1}$, and inject ranking information via trie-based marginalization. While hurting R@1 performance, Table˜2 improves R@K for all $K>1$ as well as the aggregate nDCG score. Trie-based marginalization can thus be viewed as an economical alternative to the item-level reweighting, even if it performs second to item-level reweighting under the assumption of unlimited training examples. Additionally, the improved ESCI ranking scores highlight the synergy between this token-level loss and the ESCI docIDs, which we specifically learned to be trie-friendly. We hope this result stimulates future research into optimal combinations of docID learning algorithms with token-based ranking losses.
5.3 Comparison with DEs and CEs
We compare (pointwise) autoregressive ranking with DEs and CEs. We conduct our experiment on WordNet, and report the results in Figure˜3(a). It can be seen that autoregressive ranking is on par with CEs while being significantly more powerful than DEs. Below we provide details on our choice of DEs and CEs, and describe the hyperparameter settings in Section˜B.2 and Section˜B.3, respectively.
[…]
where $\tau$ is a scalar that we fix to be 0.05. Note that this formulation incorporates the reweighting factor $\lambda(r)$. If $\lambda(r)=1$ then the loss reduces to the standard batch softmax loss commonly used in information retrieval (wu2024effectiveness). Here, we follow the same idea as in (3) and use $\lambda(r)=\frac{1}{r^{\alpha}}$. We report the results for varying $\alpha$ in Figure˜3(b), which shows that an increasing $\alpha$ improves Recall@k@k, particularly for small $k$. This result shows that our reweighting factor can be useful broadly for ranking problems, beyond the particular autoregressive ranking approach.
For CE, we concatenate the embeddings into a $2n$-dimensional vector which is passed to an MLP $f_{\text{MLP}}(\operatorname{concat}(\operatorname{E}_{\phi}(q),\operatorname{E}_{\phi}(d));\psi):\mathbb{R}^{2n}\mapsto\mathbb{R}$, where $\psi$ is trainable parameters. We use ReLU activation function, three hidden activations of size $2n$, and a final output of dimension $1$ as the similarity. To train a CE, one cannot use the same batch softmax loss as for the DEs since scoring each query against all in-batch negatives is costly. Here, for a batch of positive data $\mathcal{P}=\{(q_{i},d_{i},r_{i})\}_{i=1}^{B}$ we create a negative set of the same size given by $\mathcal{N}=\{(q_{i},d_{\rho(i)})\}_{i=1}^{B}$, where $\rho:\{1,\ldots,B\}\mapsto\{1,\ldots,B\}$ is a is a randomly-chosen permutation. Then we minimize:
$\begin{split}\mathcal{L}_{\text{CE}}(\mathcal{P},\mathcal{N};\phi,\psi)=-\sum_{i=1}^{B}\Bigg(&\frac{\lambda(r_{i})}{\sum_{j=1}^{B}\lambda(r_{j})}\log\sigma\Big(f_{\text{MLP}}(\operatorname{concat}(\operatorname{E}_{\phi}(q_{i}),\operatorname{E}_{\phi}(d_{i}));\psi)\Big)\\ &+\frac{1}{B}\log\sigma\Big(-f_{\text{MLP}}(\operatorname{concat}(\operatorname{E}_{\phi}(q_{i}),\operatorname{E}_{\phi}(d_{\rho(i)}));\psi)\Big)\Bigg),\end{split}$ (6)
with $\sigma(\cdot)$ the sigmoid function. A reweighting term $\lambda(r_{i})$ is again added to positive pairs (effect illustrated in Figure˜3(c)).
6 Conclusion
In this paper, we established a theoretical foundation for Autoregressive Ranking, proving that while DEs require embedding dimensions to grow with corpus size, ARR models generating multi-token docIDs can solve complete ranking tasks with a constant hidden dimension, given a mild condition on the rank of the embedding matrix for the docID tokens. We proposed a generalized rank-aware training loss for (pointwise) autoregressive ranking that relies on item-level reweighting and prefix-tree marginalization to distribute probability mass over valid docID tokens based on their ground-truth relevance. Experiments on WordNet and ESCI show this approach successfully suppresses invalid docID generations and improves on key ranking metrics.
7 Acknowledgements
Appendix A Proofs
A.1 Proof of Section˜3.2
[…]
@@ Appendix A Proofs
Recall from Equation˜9 that $\sum_{i\in\mathcal{P}}h_{i}z^{\prime}_{i}=\sum_{i\in\mathcal{N}}|h_{i}|z^{\prime}_{i}$. Now,
@@
Since $aS<bS$, we get $\sum_{i\in\mathcal{P}}h_{i}z^{\prime}_{i}<\sum_{i\in\mathcal{N}}|h_{i}|z^{\prime}_{i}$, which is a contradiction ($\bot$) with Equation˜9.
@@
A.2 Proof of Corollary˜3.1
[…]
@@
($\Rightarrow$) The proof is entirely analogous to the forward proof of Section˜3.2, with the exception that the contradiction targets permutations in which all token indices for which $h$ takes negative values precede all token indices for which $h$ takes negative values.